OpenStudy (anonymous):

A soup company wants to manufacture a can in the shape of a right circular cylinder that will hold 500 cm^3 of liquid. The material for the top and bottom costs 0.02 cent/cm^2, and the material for the sides costs 0.01 cent/cm^2. a) Estimate the radius r and height h of the can that costs the least to manufacture. [suggestion: Express the cost C in terms of r.]

OpenStudy (anonymous):

first find an expression the area of the can in terms of r and h hint: the area of the can is the area of two circles of radius r and one rectangle (think about 'unrolling' the side of the can to get an expression for this area)

OpenStudy (anonymous):

So far I have \[500 = \pi r ^{2}h\] and SA=\[2 \pi r ^{2}+2 \pi rh\]

OpenStudy (anonymous):

Good...then find an expression for the cost of the can: area of ends (the two circles) times .02 plus area of the side times .01

OpenStudy (anonymous):

I don't understand how to do that part

OpenStudy (anonymous):

That's the equation you're going to want to minimize (ie find the derivative and set it equal to zero)

OpenStudy (anonymous):

area of the ends is 2*pi*r^2 so cost of material used on the ends is .02*2*pi*r^2

OpenStudy (anonymous):

so Cost = .02*2*pi*r^2 +.01*2*pi*r*h

OpenStudy (anonymous):

That part clear?

OpenStudy (anonymous):

yes. now what do I do with those equations?

OpenStudy (anonymous):

like I said, need to minimize cost.. so take the derivative and set equal to zero...

OpenStudy (anonymous):

one little problem though.. it's an equation in two variables...

OpenStudy (anonymous):

we'd like it to be in terms of only one variable (r)...

OpenStudy (anonymous):

so... is there more info in the problem statement that you can use to get a second equation containing r and h?

OpenStudy (anonymous):

there is :)

OpenStudy (anonymous):

they tell you the volume of the can...

OpenStudy (anonymous):

find that equation, then you'll see that you can use it to simplify the equation we found for Cost into an equation with only one variable (r)

OpenStudy (anonymous):

so the 500=pi*r^2*h?

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

solve for 'h' and sub.s into the cost equation we found...

OpenStudy (anonymous):

500/(pr*r^2) = h right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay now what do I do after I plug that into the cost equation?

OpenStudy (anonymous):

you're in first year calc. right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, it's a minimization problem... you take the derivative of cost (with respect to radius) and set it equal to zero, remember? ... that tells you the equation for 'r' that gives you a max or min for cost...

OpenStudy (anonymous):

we know here that when Cost' =0 we're going to have a minimum... so don't be concerned about determining whether Cost'=0 is a max...

OpenStudy (anonymous):

I'm confused again. I don't think I'm supposed to be dealing with derivatives in this section and I forgot how to find the derivative of this function anyways....

OpenStudy (anonymous):

Is the chapter named 'optimization' ?

OpenStudy (anonymous):

No, it's just functions

OpenStudy (anonymous):

the chapter is named 'functions' ??

OpenStudy (anonymous):

yes and the section I'm working on is 'functions and the analysis of graphical information'

OpenStudy (anonymous):

I think they probably expect you to take the derivative odf cost and set it equal to zero and then solve for r.... if not then there are other ways ... you could plot the cost function on your calculator or some software and trace until you are near the min...

OpenStudy (anonymous):

yes then that's probably what they want you to do...

OpenStudy (anonymous):

plot the cost function we found Cost(r) ... trace it to locate the minimum cost... find the r value at that minimum point

OpenStudy (anonymous):

I'm just getting a line when I plug it into my calculator

OpenStudy (anonymous):

use 'x' for 'r'

OpenStudy (anonymous):

I did

OpenStudy (anonymous):

it's not a line... it has a squared term and a 1/r term in it... make sure you entered it correctly... make sure the scale is zoomed in enough so that you can see features...

OpenStudy (anonymous):

I plugged it in as \[(.02)2 \pi x ^{2} + (.01)2 \pi x(500/ \pi x ^{2})\]

OpenStudy (anonymous):

sure.. you can simplify it a bit before you enter it... that might help

OpenStudy (anonymous):

anyway the x window should be set to around 0...6

OpenStudy (anonymous):

you'll see a definite minimum then with the y window set to 0..6 as well

OpenStudy (anonymous):

 I have something like this now